A triangular prism is 18.1 centimeters long and has a triangular face with a base of 16 centimeters and a height of 15 centimeters. The other two sides of the triangle are each 17 centimeters. What is the surface area of the triangular prism?
Answer by rothauserc(4718)

surface area(SA) = 2 * (area of the base) + (perimeter of the base) * L where area of the base is (1/2)bh, L is length, and perimeter of base is S1 + S2 + S3
SA = 2*(1/2)*16*15 + (16+17+17)*18.1
SA = 240 + 905
SA = 1145 square centimeters